OpenStudy (anonymous):

–6x + 5y ≥ –30 –6x + 5y ≤ –30 5x – 6y ≤ –30 5x – 6y ≥ –30

OpenStudy (anonymous):

can someone show me how to make these in y=mx+b form

OpenStudy (anonymous):

you just have to get y by itself on each one. like on the first one: -6x+5y=-30 5y=6x-30 then divide by five on both sides y=6/5x-6 I really hope this helps :) good luck, and if u need more just ask :)

OpenStudy (anonymous):

does the negative go away for each one @nickirivera

OpenStudy (anonymous):

no, I added 6x to both sides so that it eliminates on the lleft and so that the x is on the right for slope intercept form

OpenStudy (anonymous):

o kk